OpenStudy (anonymous):

A ship sails 20km roughly North-East, it then turns and sails 12km Due west until it is directly North of it’s starting position. How far North of where it started is it?

OpenStudy (anonymous):

400 - 144 = 256

OpenStudy (sriram):

16km

OpenStudy (anonymous):

Pythagorean theorem again  x^2+12^2=20^2 x^2+144=400 x^2=400-144 x=sqrt(400-144) x=16